Christina makes a $9.25 purchase at a bookstore in Reno with a $20 bill. The store has no bills and gives her the change in quarters and 50-cent piece. There are 30 coins in all. How many of each kind are there?
Answer by stanbon(75887)

Christina makes a $9.25 purchase at a bookstore in Reno with a $20 bill. The store has no bills and gives her the change in quarters and 50-cent piece. There are 30 coins in all. How many of each kind are there?
--------------------------
Change = 20 - 9.25 = $10.75
-----------------------------
Equations:
Quantity Equation: q + f = 30
Value Equation:..25q +50f= 1075
-------------------------------------
Modify to set up for elimination:
25q + 25f = 25*30
25q + 50f = 1075
----------------------
Subtract 1st equation from the 2nd to solve for "f":
25f = 325
f = 13 (# of fifty-cent pieces)
---
Substitute to solve for "q":
q + 13 = 30
q = 17 (# of quarters)
==================================
Cheers,
Stan H.